OpenStudy (anonymous):

Find the least common denominator (LCD) for the fractions. 1/12 and 3/4

OpenStudy (anonymous):

@IGREEN. @Data_LG2

OpenStudy (anonymous):

What's the smallest number 12 and 4 can get into?

OpenStudy (anonymous):

ill try but i havent done it for a while

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

or do you mean by divison

OpenStudy (anonymous):

No, 1 / 4 = 0.25 We need whole numbers.

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

Err..no..we want a number that both 4 and 12 can multiply to get.

OpenStudy (anonymous):

4 * 2 = 8 8 / 12 = 0.67 12 cannot get into 8, so 8 is incorrect.

OpenStudy (anonymous):

4*9=36 and 36/12=3?

OpenStudy (anonymous):

Yes, good, that's how you find them..but that's too big..there's a smaller number.

OpenStudy (anonymous):

What's 4 * 3?

OpenStudy (anonymous):

this is hard. um 4*6=24 and 24/12=2

OpenStudy (anonymous):

4*3=12

OpenStudy (anonymous):

12/12=1

OpenStudy (anonymous):

Yes, you got it.

OpenStudy (anonymous):

So 12 is the LCD here..you always want to first check and see if one of the denominators can get into the other.